What are the chances of rolling at least one six in four rolls of a die?
About 52%
The answer
P(at least one 6) = 1 - (5/6)^4 = 1 - 625/1296 ≈ 51.8%.
Source: Verified
- logic-riddles
More from Logic & Riddles
- A bat and ball cost $1.10. The bat costs $1 more than the ball. What does the ball cost?
- If 5 machines take 5 minutes to make 5 widgets, how long for 100 machines to make 100 widgets?
- A lily pad doubles in size every day. It covers the lake in 48 days. When was it half covered?
- You have two ropes that each take 1 hour to burn (unevenly). How do you measure 45 minutes?
- A farmer has 17 sheep. All but 9 die. How many are left?
- Two fathers and two sons go fishing. They catch 3 fish, one each. How?